# One-Layer Transformers are Provably Optimal for In-context Reasoning and Distributional Association Learning in Next-Token Prediction Tasks

### v2

@@ A.4 Directional convergence of running gradient descent on the joint query-key matrix
Appendix B Missing Proofs in Section 4
B.1 Proof of Lemma 4.1
Lemma B.1.
@@ Appendix B Missing Proofs in Section 4
(Optimality of linear attention for noisy task) Under Assumption 2.2 and the reparameterization regime defined in Lemma 4.1, for all $\alpha\in(0,1)$, using linear attention we obtain
@@
It suffices to show that
@@
$L(\lambda,\gamma)$ $=\mathbb{E}_{y,z_{1:H+1}}\left[-\ln\frac{\exp(\xi_{z_{H+1}})}{\sum_{j\in[N+1]}\exp(\xi_{j})}\right]=\mathbb{E}_{y}\left[\mathbb{E}_{z_{1:H+1}}\left[-\ln\frac{\exp(\xi_{z_{H+1}})}{\sum_{j\in[N+1]}\exp(\xi_{j})}\mid y\right]\right]$ (25)
$=\mathbb{E}_{y}\left[\mathbb{E}_{z_{1:H}}\left[(\alpha-1)\left(\ln\frac{e^{C_{q,y}\lambda}}{e^{C_{q,y}\lambda}+e^{\beta}+N-1}\right)-\alpha\left(\ln\frac{e^{\beta}}{e^{C_{q,y}\lambda}+e^{\beta}+N-1}\right)\right]\right].$
@@
for any bounded $0\leq C\leq H$. Note that we require $\alpha$ strictly larger than 0 so that we can use $\lim_{\lambda\to\infty}\ln(\frac{e^{\beta}}{e^{C_{q,y}\lambda}+e^{\beta}+N-1})=\ln(\lim_{\lambda\to\infty}\frac{e^{\beta}}{e^{C_{q,y}\lambda}+e^{\beta}+N-1})=\ln\alpha$.
Fix an output token $y$. Similar to the proof of Lemma 3.1, we start by examining the attention scores ${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$ for $h=1,2,\dots,H$.
  • For $z_{h}=\tau$, we have
    ${\bm{W}}{\bm{x}}_{h}$ $=\lambda\left(E(q)\tilde{E}^{\top}(q)-E(q)E^{\top}(\tau)\right)(E(\tau)+\tilde{E}(z_{h-1}))$ (26)
    $=\lambda\left(\mathbb{1}\{z_{h-1}=q\}-1\right)E(q)=\mathbf{0}.$
@@
    ${\bm{W}}{\bm{x}}_{h}$ $=\lambda\left(E(q)\tilde{E}^{\top}(q)-E(q)E^{\top}(\tau)\right)(E(z_{h})+\tilde{E}(z_{h-1}))$ (27)
    $=\lambda\mathbb{1}\{z_{h-1}=q\}E(q).$
@@
Next, we compute $\xi_{F,j}$ for $j\in[N]$. We have ${\bm{F}}=E(\tau)(\gamma E^{\top}(q)+\tilde{E}^{\top}(q))$ and
$\xi_{F}$ $={\bm{U}}{\bm{F}}\left({\bm{x}}_{H}+\sum_{h=1}^{H}({\bm{x}}_{H}^{\top}W{\bm{x}}_{h}){\bm{x}}_{h}\right)$
$={\bm{U}}E(\tau)(\gamma E^{\top}(q)+\tilde{E}^{\top}(q))\left({\bm{x}}_{H}+\sum_{h=1}^{H}({\bm{x}}_{H}^{\top}W{\bm{x}}_{h}){\bm{x}}_{h}\right)$
$={\bm{e}}_{\tau}(\gamma E^{\top}(q)+\tilde{E}^{\top}(q))(E(q)+\tilde{E}(z_{H-1})+\lambda C_{q,y}(E(y)+\tilde{E}(q))=(\gamma+\lambda C_{q,y}){\bm{e}}_{\tau}.$

### v3

@@ B.4 Directional convergence of running gradient descent on the joint query-key matrix
Appendix C Missing Proofs in Section 4
C.1 Proof of Lemma 4.1
Proof.
Consider a sentence with a trigger $q$ and output $y$. Similar to the proof of Lemma 3.1, we first compute the pre-softmax attention scores ${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$. We have
${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$ $=(E(q)+\tilde{E}(z_{h-1}))^{\top}\left(\sum_{q^{\prime}\in{\mathcal{Q}}}\lambda_{q^{\prime}}E(q^{\prime})\left(\tilde{E}(q^{\prime})^{\top}-\sum_{x=1,x\neq q^{\prime}}^{N}\tilde{E}(x)^{\top}\right)\right){\bm{x}}_{h}$ (19)
$=\lambda_{q}\left(\tilde{E}(q)^{\top}-\sum_{x=1,x\neq q}^{N}\tilde{E}(x)^{\top}\right)(E(z_{h})+\tilde{E}(z_{h-1}))$ (20)
$=\lambda_{q}\left(\tilde{E}(q)^{\top}-\sum_{x=1,x\neq q}^{N}\tilde{E}(x)^{\top}\right)\tilde{E}(z_{h-1})$ (21)
$=\lambda_{q}\left(\mathbb{1}\{z_{h-1}=q\}-\mathbb{1}\{z_{h-1}\neq q\}\right).$ (22)
It follows that
${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}=\begin{cases}\lambda_{q}&\qquad\text{if }z_{h-1}=q,\\ -\lambda_{q}&\qquad\text{otherwise}.\end{cases}$ (23)
The attention score at the $h$-th token in a sentence is
$\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})$ $=\frac{\exp({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})}{\sum_{j=1}^{H}\exp({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{j})}=\frac{\exp(\lambda_{q}(\mathbb{1}\{z_{h-1}=q\}-\mathbb{1}\{z_{h-1}\neq q\}))}{\sum_{j=1}^{H}\exp(\lambda_{q}(\mathbb{1}\{z_{h-1}=q\}-\mathbb{1}\{z_{h-1}\neq q\}))}$ (24)
$=\begin{cases}&\frac{\exp(\lambda_{q})}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda_{q})}\quad\text{if }z_{h-1}=q,\\ &\frac{\exp(-\lambda_{q})}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda_{q})}\quad\text{otherwise.}\end{cases}$ (25)
Obviously, $\lim_{\lambda_{q}\to\infty}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})=1$ if $z_{h-1}=q$ and $\lim_{\lambda_{q}\to\infty}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})=0$ otherwise.
Next, we compute $\xi_{j}$ for $j=y$ and $j\neq y$. Recall that ${\bm{V}}=s{\bm{I}}_{d}$ and ${\bm{e}}_{y}^{\top}{\bm{U}}{\bm{x}}_{h}=\mathbb{1}\{z_{h}=y\}$. With $j=y$, we have
$\xi_{y}$ $={\bm{e}}_{y}^{\top}{\bm{U}}{\bm{V}}\sum_{h=1}^{H}\sigma({\bm{x}}_{H}{\bm{W}}{\bm{x}}_{h}){\bm{x}}_{h}=s\sum_{h=1}^{H}\sigma({\bm{x}}_{H}{\bm{W}}{\bm{x}}_{h})\mathbb{1}\{z_{h}=y\}$ (26)
$=s(\sum_{z_{h-1}=q}\sigma({\bm{x}}_{H}{\bm{W}}{\bm{x}}_{h})\mathbb{1}\{z_{h}=y\}+\sum_{z_{h-1}\neq q}\sigma({\bm{x}}_{H}{\bm{W}}{\bm{x}}_{h})\mathbb{1}\{z_{h}=y\})$ (27)
$=s\frac{C_{q,y}\exp(\lambda_{q})+(C_{y}-C_{q,y})\exp(-\lambda_{q})}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda_{q})}.$ (28)
With $j\neq y$, we have
$\xi_{j}$ $={\bm{e}}_{j}^{\top}{\bm{U}}{\bm{V}}\sum_{h=1}^{H}\sigma({\bm{x}}_{H}{\bm{W}}{\bm{x}}_{h}){\bm{x}}_{h}=s\sum_{h=1}^{H}\sigma({\bm{x}}_{H}{\bm{W}}{\bm{x}}_{h})\mathbb{1}\{z_{h}=j\}$ (29)
$=s\sum_{h=1}^{H}\frac{\exp(-\lambda_{q})}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda_{q})}\mathbb{1}\{z_{h}=j\}$ (30)
$=s\frac{C_{j}\exp(-\lambda_{q})}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda_{q})}.$ (31)
The desired statement follows from the fact that $1\leq C_{q,y}<H,0\leq C_{j}<H$ and thus $\lim_{\lambda_{q}\to\infty}\frac{C_{q,y}\exp(\lambda_{q})+(C_{y}-C_{q,y})\exp(-\lambda_{q})}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda_{q})}=1$ and $\lim_{\lambda_{q}\to\infty}\frac{C_{j}\exp(-\lambda_{q})}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda_{q})}=0$ for $j\neq y$. Hence,
$\lim_{s\to\infty}\lim_{\lambda_{q}\to\infty}\xi_{y}=\lim_{s\to\infty}s=\infty$ (32)
$\lim_{s\to\infty}\lim_{\lambda_{q}\to\infty}\xi_{j}=\lim_{s\to\infty}0=0.$ (33)
C.2 Proof of Theorem 4.2
Proof.
Consider a sample sentence with trigger word $q$ and output word $y$. We use short-hand notations $A=C_{q,y},B=C_{y},x=\lambda$. Note that $1\leq A<B<H$. The loss incurred by this sample is
$f(s,x)$ $=-\ln\frac{\exp(\xi_{y})}{\exp(\xi_{y})+\sum_{i\neq x}\exp(\xi_{i})}$ (34)
$=-\xi_{y}+\ln(\exp(\xi_{y})+\sum_{i\neq y}\exp(\xi_{i}))$
$=-s\frac{C_{q,y}\exp(x)+(C_{y}-C_{q,y})\exp(-x)}{C_{q,y}\exp(x)+(H-C_{q,y})\exp(-x)}$
$+\ln(\exp(s\frac{C_{q,y}\exp(x)+(C_{y}-C_{q,y})\exp(-x)}{C_{q,y}\exp(x)+(H-C_{q,y})\exp(-x)})+\sum_{i\neq y}\exp(s\frac{C_{i}\exp(-x)}{C_{q,y}\exp(x)+(H-C_{q,y})\exp(-x)}))$
$=-s\frac{A\exp(x)+(B-A)\exp(-x)}{A\exp(x)+(H-A)\exp(-x)}$
$+\ln(\exp(s\frac{A\exp(x)+(B-A)\exp(-x)}{A\exp(x)+(H-A)\exp(-x)})+\sum_{i\neq y}\exp(s\frac{C_{i}\exp(-x)}{A\exp(x)+(H-A)\exp(-x)}))$
$=-s\frac{Ae^{2x}+B-A}{Ae^{2x}+H-A}+\ln(\exp(s\frac{Ae^{2x}+B-A}{Ae^{2x}+H-A})+\sum_{i\neq y}\exp(s\frac{C_{i}}{Ae^{2x}+H-A})).$
Let $g(x)=Ae^{2x}+H-A$ and $u(x)=Ae^{2x}+B-A$. We have
$f(s,x)=-s\frac{u(x)}{g(x)}+\ln(\exp(s\frac{u(x)}{g(x)})+\sum_{i\neq y}\exp(s\frac{C_{i}}{g(x)})).$ (35)
Let $v(s,x)=\exp(s\frac{u(x)}{g(x)})+\sum_{i\neq y}\exp(s\frac{C_{i}}{g(x)})$. Note that $v(s,x)>0$ for all $s,x\in\mathbb{R}$.
Next, we compute the partial derivatives of $f$ with respect to $x$ and $s$. We have
$\frac{dg}{dx}$ $=2Ae^{2x}$ (36)
$\frac{du}{dx}$ $=2Ae^{2x}$ (37)
$\frac{d\frac{u}{g}}{dx}$ $=\frac{u^{\prime}(x)g(x)-u(x)g^{\prime}(x)}{g(x)^{2}}$ (38)
$=\frac{2Ae^{2x}(Ae^{2x}+H-A)-2Ae^{2x}(Ae^{2x}+B-A)}{g(x)^{2}}$ (39)
$=\frac{2Ae^{2x}(H-B)}{g(x)^{2}}$ (40)
$\frac{d\frac{1}{g}}{dx}$ $=-\frac{g^{\prime}(x)}{g(x)^{2}}=\frac{-2Ae^{2x}}{g(x)^{2}}.$ (41)
Additionally,
$\frac{\partial v}{\partial x}$ $=s\frac{d\frac{u}{g}}{dx}\exp(s\frac{u(x)}{g(x)})+\sum_{i\neq y}sC_{i}\frac{d\frac{1}{g}}{dx}\exp(s\frac{C_{i}}{g(x)})$ (42)
$=s\left(\frac{2Ae^{2x}(H-B)}{g(x)^{2}}\exp(s\frac{u(x)}{g(x)})+\sum_{i\neq y}-2C_{i}\frac{Ae^{2x}}{g(x)^{2}}\exp(s\frac{C_{i}}{g(x)})\right)$ (43)
$=\frac{2Ase^{2x}}{g(x)^{2}}\left((H-B)\exp(s\frac{u(x)}{g(x)})-\sum_{i\neq y}C_{i}\exp(s\frac{C_{i}}{g(x)})\right),$ (44)
and
$\frac{\partial v}{\partial s}$ $=\frac{u(x)}{g(x)}\exp(s\frac{u(x)}{g(x)})+\sum_{i\neq y}\frac{C_{i}}{g(x)}\exp(s\frac{C_{i}}{g(x)})$ (45)
$=\frac{1}{g(x)}\left(u(x)\exp(s\frac{u(x)}{g(x)})+\sum_{i\neq y}C_{i}\exp(s\frac{C_{i}}{g(x)})\right)$ (46)
It follows that
$\frac{\partial f}{\partial x}$ $=-s\frac{d\frac{u}{g}}{dx}+\frac{\frac{\partial v}{\partial x}}{v(x)}$ (47)
$=-\frac{2Ase^{2x}(H-B)}{g(x)^{2}}+\frac{1}{v(x)}\frac{2Ase^{2x}}{g(x)^{2}}\left((H-B)\exp(s\frac{u(x)}{g(x)})-\sum_{i\neq y}C_{i}\exp(s\frac{C_{i}}{g(x)})\right)$ (48)
$=-\frac{2Ase^{2x}}{g(x)^{2}v(x)}\left(v(x)(H-B)-(H-B)\exp(s\frac{u(x)}{g(x)})+\sum_{i\neq y}C_{i}\exp(s\frac{C_{i}}{g(x)})\right)$ (49)
$=-\frac{2Ase^{2x}}{g(x)^{2}v(x)}\left(\sum_{i\neq y}(H-B+C_{i})\exp(s\frac{C_{i}}{g(x)})\right).$ (50)
Since $A>0,H-B+C_{i}>0$ and $v(x)>0$, we have $\frac{\partial f}{\partial x}<0$ whenever $s>0$.
Next, we have
$\frac{\partial f}{\partial s}$ $=-\frac{u(x)}{g(x)}+\frac{\frac{\partial v}{\partial s}}{v(x)}$ (51)
$=\frac{1}{g(x)v(x)}\left(-u(x)v(x)+u(x)\exp(s\frac{u(x)}{g(x)})+\sum_{i\neq y}C_{i}\exp(s\frac{C_{i}}{g(x)})\right)$ (52)
$=\frac{1}{g(x)v(x)}\left(\sum_{i\neq y}(C_{i}-u(x))\exp(s\frac{C_{i}}{g(x)})\right)$ (53)
$=-\frac{1}{g(x)v(x)}\left(\sum_{i\neq y}(Ae^{2x}+B-A-C_{i})\exp(s\frac{C_{i}}{g(x)})\right).$ (54)
[…]
  • In the first phase $1\leq t\leq T_{0}$, we first show that $s_{t}$ may fluctuate but is always positive. Recall that for scalar value of $s$, the normalize gradient descent is equal to sign descent $s_{t}=s_{t-1}-\eta\operatorname{sign}{\frac{\partial L}{\partial s}}$. In the worst case, the signs of the partial derivatives are always positive. It follows that
[…]
    Thus, $s_{t}>0$ always holds in the first phase. This implies that $\frac{\partial f}{\partial\lambda_{q}}<0$ in the first phase. Therefore, the update formula $|\lambda_{q,t}=\frac{\eta t}{\absolutevalue{{\mathcal{Q}}}}$ always holds, which implies
[…]
    This implies that $\frac{\partial L}{\partial s}<0$.
  • In the second phase $t>T_{0}$, we now have that the signs of all partial derivatives are negative. Therefore,
    $s_{t}$ $\geq\frac{1}{2}+\eta(t-T_{0}),$ (57)
    $\lambda_{q,t}$ $|.\displaystyle=\frac{\eta t}{\absolutevalue{{\mathcal{Q}}}}$. (58)
    Plugging these into (35), we obtain
    $f(s_{t},\lambda_{q,t})$ $=\ln(1+\sum_{i\neq y}\frac{\exp(s_{t}\frac{C_{i}}{g(\lambda_{q,t})})}{\exp(s_{t}\frac{u(\lambda_{q,t})}{g(\lambda_{q,t})})})=\sum_{i\neq y}\frac{\exp(s_{t}\frac{C_{i}}{g(\lambda_{q,t})})}{\exp(s_{t}\frac{u(\lambda_{q,t})}{g(\lambda_{q,t})})}$ (59)
    $\leq N\frac{\exp(s_{t}\frac{H}{g(\lambda_{q,t})})}{\exp(s_{t}\frac{u(\lambda_{q,t})}{g(\lambda_{q,t})})}\leq O\left(N\frac{\exp(s_{t}\frac{H}{g(\lambda_{q,t})})}{\exp(2s_{t})}\right)$ (60)
    $\leq O(N\exp(-\eta t)),$ (61)
    where the second inequality is from $\frac{u(x)}{g(x)}\geq\frac{1}{2}$ for sufficiently large $x$, and the last inequality is from $2s_{t}=\Omega(\eta t)$ and
    $s_{t}\frac{H}{g(\lambda_{q,t})}$ $\leq(s_{0}+\eta T_{0}+\eta t)\frac{H}{C_{q,y}e^{2\lambda_{q,t}}+H-C_{q,y}}$ (62)
[…]
    $\leq O(1).$ (64)
Appendix D Missing Proofs in Section 5
D.1 Proof of Lemma 5.1
We prove the following two lemmas on the optimality of linear and softmax attention for the noisy setting.
Lemma D.1.
@@ Appendix D Missing Proofs in Section 5
(Optimality of linear and ReLU attention for noisy task) Under the reparameterization regime defined in Lemma 5.1, for all $\alpha\in(0,1)$, using linear and ReLU attention, we obtain
@@
Fix a sentence with trigger $q$ and output $y$. It suffices to show that
@@
$L(\lambda,\gamma)$ $=\mathbb{E}_{q,y,z_{1:H+1}}\left[-\ln\frac{\exp(\xi_{z_{H+1}})}{\sum_{j\in[N+1]}\exp(\xi_{j})}\right]=\mathbb{E}_{y}\left[\mathbb{E}_{z_{1:H+1}}\left[-\ln\frac{\exp(\xi_{z_{H+1}})}{\sum_{j\in[N+1]}\exp(\xi_{j})}\mid y\right]\right]$ (67)
$=\mathbb{E}_{q,y}\left[\mathbb{E}_{z_{1:H}}\left[(\alpha-1)\left(\ln\frac{e^{C_{q,y}\lambda}}{e^{C_{q,y}\lambda}+e^{C_{q,y}\lambda+\gamma}+N-1}\right)-\alpha\left(\ln\frac{e^{C_{q,y}\lambda+\gamma}}{e^{C_{q,y}\lambda}+e^{C_{q,y}\lambda+\gamma}+N-1}\right)\right]\right].$
@@
for any bounded $0\leq C\leq H$. Note that we require $\alpha$ strictly larger than 0 so that we can use $\lim_{\lambda\to\infty}\ln(\frac{e^{C_{q,y}\lambda+\gamma}}{e^{C_{q,y}\lambda}+e^{C_{q,y}\lambda+\gamma}+N-1})=\ln(\lim_{\lambda\to\infty}\frac{e^{C_{q,y}\lambda+\gamma}}{e^{C_{q,y}\lambda}+e^{C_{q,y}\lambda+\gamma}+N-1})=\ln\alpha$.
We turn to proving (66). Similar to the proof of Lemma 3.1, we start by examining the attention scores ${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$ for $h=1,2,\dots,H$. First, the product ${\bm{x}}_{H}^{\top}{\bm{W}}$ is equal to
$(E(q)^{\top}+\tilde{E}(z_{H-1})^{\top})\lambda\left(\sum_{q^{\prime}\in{\mathcal{Q}}}E(q^{\prime})\left(\tilde{E}^{\top}(q^{\prime})-E^{\top}(\tau)\right)\right)$ (68)
$=\lambda E(q)^{\top}\left(\sum_{q^{\prime}\in{\mathcal{Q}}}E(q^{\prime})\left(\tilde{E}^{\top}(q^{\prime})-E^{\top}(\tau)\right)\right)$ (69)
$=\lambda(\tilde{E}(q)^{\top}-E(\tau)^{\top}).$ (70)
Next, we consider two cases:
  • For $z_{h}=\tau$, we have $z_{h-1}=q$, therefore
    ${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$ $=\lambda\left(\tilde{E}^{\top}(q)-E^{\top}(\tau)\right)(E(\tau)+\tilde{E}(q))=\mathbf{0}.$ (71)
@@
    ${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$ $=\lambda\left(\tilde{E}^{\top}(q)-E^{\top}(\tau)\right)(E(z_{h})+\tilde{E}(z_{h-1}))=\lambda\mathbb{1}\{z_{h-1}=q\}.$ (72)
@@
Next, we compute $\xi_{F,j}$ for $j\in[N]$. We have ${\bm{V}}={\bm{I}}_{d}$, ${\bm{F}}=E(\tau)\left(\sum_{q^{\prime}\in{\mathcal{Q}}}\gamma E^{\top}(q^{\prime})+\tilde{E}^{\top}(q^{\prime})\right)$ and
$\xi_{F}$ $={\bm{U}}{\bm{F}}\left({\bm{x}}_{H}+\sum_{h=1}^{H}({\bm{x}}_{H}^{\top}W{\bm{x}}_{h}){\bm{V}}{\bm{x}}_{h}\right)$
$={\bm{U}}E(\tau)\left(\sum_{q^{\prime}\in{\mathcal{Q}}}\gamma E^{\top}(q^{\prime})+\tilde{E}^{\top}(q^{\prime})\right)\left({\bm{x}}_{H}+\sum_{h=1}^{H}({\bm{x}}_{H}^{\top}W{\bm{x}}_{h}){\bm{x}}_{h}\right)$
$={\bm{e}}_{\tau}\left(\sum_{q^{\prime}\in{\mathcal{Q}}}\gamma E^{\top}(q^{\prime})+\tilde{E}^{\top}(q^{\prime})\right)(E(q)+\tilde{E}(z_{H-1})+\lambda C_{q,y}(E(y)+\tilde{E}(q)))$
$={\bm{e}}_{\tau}\left(\gamma E^{\top}(q)+\tilde{E}^{\top}(q)\right)(E(q)+\tilde{E}(z_{H-1})+\lambda C_{q,y}(E(y)+\tilde{E}(q)))$
$=(\gamma+\lambda C_{q,y}){\bm{e}}_{\tau},$
where the last two equalities uses the fact that $z_{H-1}$ cannot be a trigger word, otherwise the condition IV in the data model 2.1 would be violated.